### v1

# Promptus: Can Prompts Streaming Replace Video Streaming with Stable Diffusion

Abstract
With the exponential growth of video traffic, traditional video streaming systems are approaching their limits in compression efficiency and communication capacity. To further reduce bitrate while maintaining quality, we propose Promptus, a disruptive novel system that streaming prompts instead of video content with Stable Diffusion, which converts video frames into a series of "prompts" for delivery. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. To the best of our knowledge, Promptus is the first attempt to replace video codecs with prompt inversion and the first to use prompt streaming instead of video streaming. Our work opens up a new paradigm for efficient video communication beyond the Shannon limit.
1 Introduction
With the rapid development of streaming applications (such as video-on-demand [2, 8], live video [14, 7], video conferencing [15, 11], cloud gaming [3, 13], etc.), the traffic of network video has been continuously growing. To reduce traffic, video codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5] are widely used to compress videos. These codecs achieve compression by removing spatial and temporal redundancies. However, these redundancies are limited, so there is an upper bound on the compression ratio (subject to the Shannon limit [40]). In order to further compress the video, non-redundant content in the video will be discarded, which will greatly degrade the video quality, such as causing blurring and blocking artifacts.footnotetext: Promptus will be released at https://pku-netvideo.github.io/ once published.
Figure 1: Promptus can invert a given image into a prompt. Based on this prompt, Stable Diffusion can generate an almost identical image to the original. In contrast, existing methods can only generate semantically similar images.
To improve compression efficiency, some deep learning-based codecs and streaming frameworks have been proposed in recent years. Neural-embedded codecs [31, 30, 19, 37, 47] replace handcrafted modules (such as motion vector estimation [45]) in traditional codec frameworks with neural networks (such as autoencoders), learning more intelligent and efficient compression capabilities through end-to-end optimization. However, due to still following the traditional encoding framework, the improvement in compression ratio is limited. Neural-enhanced streaming [50, 49, 26, 52, 56] transmits distorted low-bitrate videos and then restores high-fidelity details using post-processing algorithms such as super-resolution. However, these post-processing algorithms rely on prior knowledge learned from training sets. Due to the domain gap, their performance will degrade in unseen scenes [48]. Generative streaming [24, 44, 42, 29] transmits small-sized semantic information, and the receiver can generate video based on these semantics. However, most generative algorithms are designed for specific tasks (e.g., talking video generation driven by facial keypoints [24, 44]), limiting their generality.
With the popularity of AIGC (AI-generated content), Stable Diffusion [38, 10, 9] has attracted extensive attention thanks to its powerful text-to-image generation capability. By pre-training on an internet-scale dataset [6], Stable Diffusion learns prior knowledge of all human visual domains, and simultaneously learns the mapping from text to images. Therefore, Stable Diffusion can generate high-fidelity images based on a brief prompt composed of a few words.
We are motivated by this question: is it possible for Stable Diffusion to replace the video codecs? During streaming, the sender streams prompts instead of streaming encoded videos and the receiver generates videos instead of decoding videos. In this way, the traffic of network video is reduced from the video scale to the text scale, achieving communication efficiency beyond the Shannon limit.
In this paper, we propose Promptus, a system that inverts video frames into Stable Diffusion prompts, achieving ultra-low bitrate video streaming. To bring this vision to fruition, we address the following technical challenges:
First, how to ensure pixel alignment between the generated frames and the real frames. To invert a frame into a prompt, the most straightforward and powerful approach is to manually write a textual description. However, the frame generated using this description can only guarantee semantic consistency with the original frame, while the differences at the pixel level are often substantial, as shown in Figure 1. To achieve pixel alignment, Promptus proposes a gradient descent-based prompt fitting framework. Specifically, the prompt is randomly initialized and then used to generate frame. The pixel-wise loss between the generated frame and the real frame is calculated. The partial derivative of the loss with respect to the prompt will be calculated and the prompt is iteratively optimized using gradient descent. To implement this framework, Promptus employs single-step denoising instead of iterative denoising to avoid higher-order derivatives. Second, Promptus uses embeddings as prompts instead of text to avoid non-differentiability. Third, Promptus uses a noisy previous frame instead of random noise to reduce latent space distance. Fourth, Promptus combines reconstruction and perceptual loss to enhance the perceptual quality of the generated frames.
Second, how to control the bitrate of the prompt. Since Promptus uses embeddings as prompts, which are matrices with fixed dimensions, the bitrate of the prompt cannot adapt to dynamic network bandwidth. To address this, Promptus proposes a low-rank bitrate control algorithm. Specifically, Promptus integrates the inverse process of low-rank decomposition into the gradient descent, directly fitting the decomposed prompt. The rank is used to control the trade-off between the quality and bitrate of the prompt. When the rank is higher, the representational capability of the prompt is better, and it can describe more details in the frame, but the data size is also larger. Therefore, Promptus adaptively selects the prompt rank based on the currently available bandwidth.
Third, how to perform inter-frame compression on prompts. Promptus inverts each frame into an independent prompt without considering the correlation between prompts across frames. Therefore, when streaming video, Promptus needs to transmit prompts for each frame, resulting in the bitrate increasing linearly with the frame rate. To address this, Promptus adds a temporal smoothing regularization during prompt fitting, ensuring that temporally close frames are also sufficiently close in the prompt space. With this, Promptus only needs to sparsely transmit prompts for a few keyframes, while the prompts for the remaining frames can be approximated through linear interpolation of the keyframe prompts.
We evaluated Promptus on test videos from different domains and under real network traces. The results show that: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE [27, 38] and H.265 [5], respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
The contributions of this paper are summarized as follows: (1) We propose Promptus, which, to the best of our knowledge, is the first attempt to replace video codecs with prompt inversion and also the first to use prompt streaming to replace video streaming (§3). (2) We propose a gradient descent-based prompt fitting framework, achieving pixel-aligned prompt inversion for the first time (§3.1). (3) We build a video streaming system based on Promptus (§4).
2 Motivation and related work
2.1 Video codec and streaming
Traditional codecs. In network video traffic, most of the content is encoded using traditional codecs represented by VP8 [16], VP9 [33], H.264 [4] and H.265 [5]. These traditional video codecs achieve compression primarily by eliminating redundancy in the video. Specifically, the codecs exploit the correlation between adjacent pixel blocks to reduce spatial redundancy in intra-frame prediction. Besides, considering the similarity between consecutive video frames, the codecs employ inter-frame prediction techniques to eliminate temporal redundancy. For moving objects, the codecs perform motion estimation and compensation to further reduce redundancy. Although these techniques achieve efficient compression, the compression ratio has an upper limit due to the finite amount of redundancy. Further compression inevitably requires discarding some non-redundant information, resulting in a drastic decrease in video quality. For example, as shown in Figure 11, when the bitrate is extremely low, videos compressed using H.265 exhibit significant blurriness.
Neural-embedded codecs. With the development of deep learning, handcrafted modules in traditional codecs are being replaced by neural networks [31, 30, 19, 37, 47]. Compared to handcrafted modules, these embedded neural networks can learn more complex and nonlinear motion patterns and intra-frame correlations, thereby achieving lower distortion at the same bitrate. Furthermore, due to the ability to optimize end-to-end, neural-embedded codecs can be trained for tasks beyond compression, such as loss resilience. However, since neural-embedded codecs still adhere to the traditional coding framework, aiming to fully remove intra-frame or inter-frame redundancy, the improvement in compression ratio is limited.
Neural-enhanced streaming. In contrast to reducing redundancy, neural-enhanced streaming [50, 49, 26, 52, 56] actively discards most of the information (including non-redundant information) and transmits highly distorted low-bitrate videos. Then, at the receiver side, neural network-based post-processing algorithms (such as super-resolution) are used to restore high-fidelity details. Neural-enhanced streaming leverages the powerful image restoration and enhancement capabilities of neural networks, effectively recovering lost textures and details, thus ensuring video quality while achieving substantial compression. However, these post-processing algorithms rely on prior knowledge learned from training datasets, such as texture features and edge structures. Due to the complex and diverse nature of real-world video, there exists domain gaps with the training set. The performance of these algorithms often degrades in unseen scenarios [48].
Generative streaming. Compared to using neural networks for post-processing, generative streaming directly uses neural networks to generate videos to handle higher compression ratios [24, 44, 42, 29]. For example, in the context of video conferencing, Face-vid2vid [44, 24] extracts facial keypoints in real-time at the sender side for transmission. At the receiver side, it generates dynamic facial videos using the keypoints and static facial images. The bitrate of facial keypoints is much lower than that of videos, thus compressing video conferences to extremely low bitrates. Similarly, Gemino [42] transmits video streams at extremely low resolution. At the receiver side, it estimates facial motion fields from the low-resolution video, then utilizes the motion fields and high-resolution facial images to generate high-resolution facial videos. Instead of driving static images, Reparo [29] proposes a token-based generative streaming. It first uses VQGAN [20] to train a codebook of facial visual features, then maps the video to latent variables using VAE [27], and finally quantizes the latent variables into tokens according to the codebook. Since tokens are indices of the codebook, the bitrate is extremely low. In general, generative streaming can greatly compress videos, but it is often designed for specific tasks (such as video conferencing) and lacks generality.
2.2 Stable diffusion
Stable Diffusion [38, 10, 9] is the most popular open-source text-to-image generative model. It learns a denoising process from 5.85 billion image-text pairs [6], enabling it to generate high-quality images by denoising the pure noise images. Different noise images lead to different generated images. Since the noise images are randomly sampled from Gaussian distribution, the images generated by Stable Diffusion are random and uncontrollable. Therefore, to control the content of the generated images, Stable Diffusion also receives user-input prompts (in natural language) as the condition for denoising, thereby generating images that align with the semantic descriptions of the prompts. Specifically, as shown in Figure 2, $p$ represents the user-input prompt. Stable Diffusion first performs word embedding and semantic extraction using the CLIP model [36], converting discrete natural language into a continuous text embedding $c$ (an $m*n$ matrix). The text embeddings and the randomly sampled noise image $N$ are input into the denoising process of Stable Diffusion, generating the denoised $N^{\prime}$. Next, $N^{\prime}$ is input into the denoising process again for $T$ iterations to obtain the denoised $Z$. Since the denoising process occurs in the latent space, $Z$ needs to be input into the VAE Decoder [27] to generate the image $x$ in the pixel space.
Figure 2: How does Stable Diffusion generate high-quality images from text prompts. The details are elaborated in §2.2.
Despite the impressive image generation capabilities of Stable Diffusion, its potential in the field of video streaming has been overlooked. However, we believe that it can significantly enhance the performance of video streaming for the following reasons:
  • Stable Diffusion can effectively reduce the bandwidth overhead of videos. It can generate high-quality images using brief prompts. As shown in Figure 2, a prompt consisting of only a few words can enable Stable Diffusion to generate images with high resolution and complex textures. In codec-based streaming, such high-quality images require high bitrates. Instead of transmitting image, if we solely transmit prompts and generate images at the receiving end, the bitrate can be reduced from the image scale to the prompt scale, greatly improving communication efficiency.
  • Stable Diffusion is generalizable to most video domains. Trained on an internet-scale dataset [6] (5.85 billion images), Stable Diffusion has learned priors for nearly all video domains present on the internet. Consequently, it does not suffer from the domain gap issue and can generate images of any style and content. This gives Stable Diffusion the potential to become a versatile paradigm for video streaming.
  • Stable Diffusion can generate images in real-time. With ongoing advancements in Stable Diffusion, variants capable of real-time image generation have emerged. For instance, StreamDiffusion [28] can generate images at a speed of 100 FPS. This enables Stable Diffusion to meet the real-time requirements of video streaming.
However, due to the inability to precisely control image generation, Stable Diffusion cannot meet the fidelity requirements of video streaming. Specifically, Stable Diffusion only defines the generation process from prompt to image, without considering the inverse process of extracting prompts from images. The most straightforward solution is to automatically extract text descriptions for target images using Image/Video Captioning algorithms [46, 23] and use these extracted descriptions as prompts to generate images. Although the generated images can semantically align with the target images, these extracted descriptions are always too high-level, resulting in generated images lacking many details or having large structural differences with the target images. In fact, even if human intelligence is used to manually describe images, it is impossible to generate images that are pixel-aligned with the target images, as shown in Figure 1. In addition to using text as prompts, ControlNet [54] can utilize images (such as masks) as prompts, controlling Stable Diffusion to generate images that align with the contours of the prompt image. However, apart from the contours, details such as color and texture cannot be aligned. In conclusion, the frames generated by Stable Diffusion cannot faithfully reproduce the original frames at the pixel level, making them unsuitable for video streaming.
To materialize the above potential benefits, Promptus is the first to successfully invert frames into prompts while ensuring pixel-level alignment. Promptus is related to some work on Text Inversion [21, 39, 25], both of which learn prompts from images in an inverse manner. However, Text Inversion aims to learn new words that represent the appearance of specific objects from images, aligning only at the semantic level. Text Inversion has also been used for image compression [35]. But similarly, the inverse prompt is only used for semantic alignment, while pixel alignment is achieved by using low-resolution images as conditions.
3 Video to Prompt Inversion
Figure 3: Workflow of Promptus’s video to prompt inversion.
This section describes how Promptus inverts frames into prompts. The workflow is illustrated in Figure 3. First, to ensure pixel alignment of the generated frames, a gradient descent-based prompt fitting framework is proposed (§3.1). Second, to control the bitrate of each prompt, a low-rank decomposition-based compression algorithm is introduced (§3.2). Third, to perform inter-frame compression on prompts, a temporally smooth prompt space is proposed (§3.3).
3.1 Gradient Descent based Prompt Fitting
Gradient Descent Framework. To obtain the inverse prompt, the most straightforward approach is to train a neural network to map the target image to a prompt. However, on one hand, this approach makes it difficult to ensure pixel alignment (like the aforementioned Captioning algorithm [46, 23]). On the other hand, as the inverse process of Stable Diffusion, this neural network needs to learn comparable knowledge, but this is very expensive (e.g., training a Stable Diffusion will cost between 600,000 and 10 million US dollars [12]). Therefore, instead of training a new neural network, we propose fully leveraging Stable Diffusion’s knowledge to infer the prompt. To this end, we adopt gradient descent to iteratively fit the prompt, with the framework shown in Figure 3. Specifically, at the beginning, the prompt is randomly initialized. Then, Stable Diffusion generates a frame based on this prompt. Since the prompt is random, the generated frame is meaningless. Third, the pixel-wise difference between the generated frame and the target frame is calculated as the loss value. Fourth, backpropagation is used to compute the gradient of the loss value with respect to the prompt. Finally, the prompt is updated using gradient descent. The above steps are iteratively executed until the loss value is sufficiently small, and the resulting prompt can satisfy pixel-aligned generation. In the above steps, Stable Diffusion is pre-trained and frozen, so it has prior knowledge. This knowledge is gradually distilled into the prompt through gradient descent fitting.
To realize the aforementioned framework, there are several key components:
Single-step denoising to avoid higher-order derivatives. As described in §2.2, Stable Diffusion uses iterative denoising to generate images. Therefore, the prompt recursively affects the generated image. This causes the gradient of the loss value with respect to the prompt to involve the computation of higher-order derivatives (such as 20th order), which introduces prohibitive computational and memory overhead. Therefore, instead of using the traditional Stable Diffusion, we adopt SD Turbo [9], a variant that can generate frames through single-step denoising. The adoption of SD Turbo allows gradient descent to only compute the first-order derivatives of the prompt, greatly improving efficiency. Although the quality of the generated frames is slightly weaker than the traditional iterative denoising Stable Diffusion, these quality losses can be compensated for in an end-to-end manner during prompt fitting.
Employing embeddings as prompts to avoid non-differentiability. Computing the gradient of the prompt through backpropagation requires the forward computation from the prompt to the loss value to be completely differentiable. However, as shown in Figure 2, the forward computation includes the CLIP module [36], which converts text from discrete natural language to continuous embeddings. This step involves indexing and table lookup, making it non-differentiable. Gradients cannot be propagated to the text, preventing gradient descent. To address this, we discard the non-differentiable CLIP module and directly use embeddings as prompts for conditioning Stable Diffusion. In this case, the forward computation is fully differentiable, allowing gradient descent to be performed on the embeddings. In the following sections, prompt refers to the embedding rather than the text.
Using a noisy previous frame instead of random noise to reduce latent space distance. According to §2.2, in addition to the prompt, the input random noise also affects the frames generated by Stable Diffusion. With the same prompt, different input noise results in different generated frames. At a high level, the input noise can be viewed as a point in the latent space, and the denoising process of Stable Diffusion actually moves this point under the control of the prompt. Therefore, the goal of Promptus is to find the inverse prompt that can move the point represented by the noise to the point of the target image in the latent space. Since we adopt a single-step denoising Stable Diffusion, if the random noise is far from the target image in the latent space, this movement cannot be completed in a single step, making it impossible to fit the inverse prompt. As shown in Figure 4(a), using random noise as input, after the loss value converges, the generated frame still has noticeable differences from the target frame, including blurring, noise artifacts, and inconsistent details. Therefore, we need to reduce the distance between the input noise and the target image in the latent space. We observe that in a video, adjacent frames are close in the latent space. Thus, we manually add noise to the previous frame as follows:
$N^{t}=(1-\gamma)*Z^{t-1}+\gamma*N^{0}$ (1)
Where $Z^{t-1}$ is the previous frame in the latent space. $N^{0}$ is a fixed noise. $\gamma$ is a hyperparameter that controls the degree of noise addition, which we set to $0.95$ in the experiment. $N^{t}$ is used as the noise input to Stable Diffusion for denoising the current frame. As shown in Figure 4(c), compared to random noise, the noisy previous frame can reduce the distance in the latent space, resulting in a generated frame that better matches the target frame.
Combining reconstruction and perceptual loss functions. To achieve pixel-aligned supervision, the most intuitive loss function is the per-pixel reconstruction loss, such as MSE. These reconstruction losses attempt to minimize the error of each pixel, while errors in high-frequency details and edges often lead to large pixel errors. Therefore, to reduce the overall error, reconstruction losses tend to abandon the fitting of edges and details, resulting in overly smooth and blurry images, as shown in Figure 4(b). To make the generated frames sharp and clear, one approach is to use perceptual loss instead of reconstruction loss, such as LPIPS [55]. Perceptual loss is based on deep learning and can estimate the subjective quality of images as perceived by the human eye. Since the human eye is highly sensitive to image details, perceptual loss can make the generated images sharper with richer details. However, perceptual loss aims to maximize the overall subjective quality of the image without focusing on the exact consistency of each pixel, leading to misalignment between the generated and target images. Therefore, to simultaneously ensure pixel alignment and subjective quality, we combine the reconstruction and perceptual loss as the fitting loss $D$:
$D=\alpha*D_{rec}(x,x_{gt})+(1-\alpha)*D_{per}(x,x_{gt})$ (2)
where $D_{rec}$ represents the reconstruction loss, which is MSE by default. $D_{per}$ represents the perceptual loss, which is LPIPS by default. $\alpha$ is a hyperparameter that balances pixel alignment and perceptual quality. In our experiments, we set $\alpha$ to 0.8. The final result is shown in Figure 4(c). It can be observed that by jointly optimizing the reconstruction loss and the perceptual loss, the images generated by Promptus ensure pixel-level alignment while maintaining a sharp appearance, making them almost identical to the ground truth images.
Figure 4: Visualization of prompt fitting results. (a) Using random noise as input. (b) Only using MSE as the loss function. (c) Ours. (d) Ground Truth. The results demonstrate that the noisy previous frame and the perceptual loss both contribute to the visual quality.
3.2 Low-rank Decomposition based Prompt Bitrate Control
According to §3.1, Promptus uses embeddings as prompts instead of text. However, embeddings are $m*n$ matrices (e.g., $1024*77$), where each element is a high-bit floating-point number (e.g., 32-bit float type), resulting in a much higher bitrate. To reduce the bitrate of prompts, Promptus has two directions: First, dimensionality reduction decreases the number of parameters in the prompt. Second, quantization reduces the number of bits for each parameter in the prompt.
Low-rank matrix decomposition. To perform dimensionality reduction on the prompt, the most straightforward method is to first fit the complete prompt and then perform dimensionality reduction algorithms such as SVD (Singular Value Decomposition) or PCA (Principal Component Analysis) on it. However, these methods only perform dimensionality reduction based on the data distribution of the prompt, ignoring the impact of prompt degradation on the generation results of Stable Diffusion. This inevitably leads to a degradation in the quality of the generated images. Therefore, instead of performing explicit dimensionality reduction, Promptus proposes to directly fit a low-dimensional prompt end-to-end, thereby reducing the quality degradation caused by dimensionality reduction. To achieve this, Promptus integrates the inverse process of CANDECOMP/PARAFAC decomposition [22] into gradient descent fitting. Specifically, Promptus calculates the embedding $c$ as follows:
$\vspace{-6mm}c=\frac{u*v}{\sqrt{r}}$ (3)
where $u$ and $v$ are two low-rank factor matrices, and $r$ is the rank of the embedding. $u$ and $v$ compose the embedding $c$ through outer product and normalization. At this point, the embedding $c$, as an intermediate variable, is no longer fitted or stored. $u$ and $v$, as the new representation of the prompt, will be randomly initialized and fitted.
The rank $r$ determines the trade-off between bitrate and quality. Compared to the embedding size of $m*n$ (e.g., $1024*77$), the total size of $u$ and $v$ is $(m+n)*r$. Therefore, reducing $r$ significantly lowers the bitrate. However, on the other hand, when Rank $r$ is smaller, the embedding is constrained to be a low-rank matrix, resulting in weaker representational capability and inability to fit high-frequency details in the image, as shown in Figure 7. Consequently, it is necessary to dynamically select the most appropriate $r$ based on the currently available network bandwidth to trade off between bitrate and quality.
Fitting-aware quantization. Although the number of parameters in the prompt has been significantly reduced through low-rank matrix decomposition, each parameter in $u$ and $v$ is still a high-bit floating-point number (such as 32-bit float type). Therefore, to further reduce the bitrate, it is necessary to quantize $u$ and $v$, reducing the number of bits for each parameter (such as lowering it to 8 bits). The most straightforward approach is to first fit $u$ and $v$ and then perform quantization. However, this post-quantization technique inevitably leads to quality loss. To address this, Promptus incorporates quantization into the fitting process, automatically compensating for the quantization loss through end-to-end gradient descent. We tested the impact of different quantization configurations on quality. Compared to traditional post-quantization, our method can reduce the number of bits from 32 to 8 with almost no quality loss.
3.3 Prompt inter-frame compression based on temporal smoothing
Figure 5: At the receiver, Promptus approximates the unreceived frames by linear interpolation in the prompt space.
According to §3.1, Promptus fits each frame of the video as an independent prompt, without considering the correlation between prompts across frames. Therefore, during video streaming, Promptus needs to transmit a prompt for each frame, resulting in a linear increase in bitrate as the frame rate rises. However, codec-based streaming can avoid this problem through inter-frame compression. Is it possible to perform inter-frame compression on prompts as well? The most straightforward solution is to reshape the prompt of each frame into a two-dimensional matrix and encode it using a video codec. However, unlike conventional videos, we found that this reshaped prompt looks like random noise, lacking any patterns, structures, or smooth regions, causing video codecs to no longer work.
For inter-frame compression of prompts, our insight is: prompts are high-level semantics, so the prompts of continuous video frames should change continuously. If two temporally close frames are also sufficiently close in the prompt space, then the prompts of the frames between these two frames can be approximated by linear interpolation. With this, during streaming, we only need to sparsely transmit the prompts of a few frames (as keyframes), and the prompts of the remaining frames can be obtained by linear interpolation of the keyframe prompts, as illustrated in Figure 5. Since only a small portion of keyframes need to transmit prompts, inter-frame compression is achieved.
To ensure that adjacent frames are sufficiently close in the prompt space, we add temporal smoothing regularization to the embedding during fitting, as follows:
$\lambda=\left\|c^{t}-c^{t-1}\right\|_{2}$ (4)
Here, $c^{t}$ represents the embedding of the frame currently being fitted, while $c^{t-1}$ denotes the embedding of the previously fitted frame. The final loss function $L$ is as follows:
$L=\beta*D+(1-\beta)*\lambda$ (5)
Where $\beta$ is a hyperparameter used to balance the fitting loss and the temporal smoothing regularization. In our experiments, we set it to 0.2.
Figure 6: Visualization of the prompt interpolation results. When applying temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
Temporal smoothing regularization works, as shown in Figure 6. Without temporal smoothing regularization, the interpolation results suffer from severe distortions, disrupting the video’s content and structure. With temporal smoothing regularization, the interpolation results not only fully preserve the video details but also successfully approximate the motion in the video.
4 Prompt Streaming System Design and Implementation
§3 shows how Promptus inverts videos into pixel-aligned prompts. In this section, we will further describe how Promptus utilizes this Prompt Inversion technique to design and implement a streaming system. Specifically, it will be divided into two parts: the sender side and the receiver side.
4.1 The sender side
On the sender side, Prompt Inversion replaces video encoding. At a high level, Promptus performs Prompt Inversion on raw frames as target images. The obtained prompts are then streamed to the receiver side for image generation and playback. The details are as follows:
Initialization for the first frame. As stated in §3.1, to improve the quality of the generated images, Promptus uses the noisy previous frame as the input noise for Stable Diffusion. However, for the first frame of the video (the frame index starts from 1), there is no previous frame $Z^{0}$. To address this, Promptus uses the VAE’s Encoder to map the first frame itself to the latent space, obtaining $Z^{0}$. Then, noise is added to $Z^{0}$ according to Equation 1. The noisy $N^{1}$ will be used as the input noise for Prompt Inversion of the first frame. Since the purpose of using the noisy previous frame is to reduce the distance between the input noise and the target image in the latent space, the noisy current frame naturally works as well. Note that this noise $Z^{0}$ will also be sent to the receiver side along with the prompt for generating the first frame image. On the other hand, due to the absence of a previous frame, Promptus does not apply temporal smoothing regularization for the Fitting of the first frame.
Re-initialization for abrupt scene changes. When the video undergoes drastic changes, such as suddenly switching to a new scene, the content difference between the previous frame and the current frame becomes significant, and the distance in the latent space is no longer close. In this case, using the noisy previous frame as described in §3.1 will no longer work. Therefore, Promptus detects abrupt scene changes. Specifically, Promptus calculates the distance between the current frame and the previous frame in the latent space, and when this distance exceeds a threshold, it is considered an abrupt scene change. In this situation, Promptus treats the video after the scene change as a new video and performs the aforementioned first frame initialization again.
Sparse prompt streaming. As stated in §3.3, through temporal smoothing regularization, the prompts of most frames can be approximated by linear interpolation of the prompts of keyframes. Therefore, Promptus defines the keyframe interval as $K$ (the choice of $K$ is discussed in §5.2), adding one keyframe every $K$ frames. Only the prompts of keyframes will be streamed to the receiver. Note that when the aforementioned abrupt scene changes occur, the difference between frames in the latent space is significant, and the interpolation of prompts no longer works. In this case, the last frame before the scene change will also become a keyframe and be sent. The frames after the scene change will start a new count.
Low-rank based adaptive bitrate. According to §3.2, the higher the bitrate of the prompt, the higher the quality of the generated image. Therefore, in prompt streaming, it is necessary to increase the prompt bitrate as much as possible while avoiding network congestion to maximize the user experience. Promptus adopts the WebRTC [1] framework for prompt streaming, which dynamically probes the currently available bandwidth. Consequently, Promptus adaptively adjusts the rank to make the prompt bitrate close to the probed bandwidth. It is worth noting that the prompts transmitted by Promptus are not encoded. Therefore, compared to codec-based streaming, Promptus can precisely control the bitrate.
4.2 The receiver side
On the receiver side, Stable Diffusion’s prompt-to-image generation replaces video decoding. The details are as follows:
Video generation based on sparse prompts. After receiving the prompt of keyframe $i$, the receiver first performs linear interpolation between the prompts of adjacent keyframes $i-k$ and $i$ to approximate the prompts of the intermediate $K-1$ frames. Afterward, these prompts are used for Stable Diffusion’s image generation. As described in §3.1, the generated $Z$ for each frame will be noised and used as the input noise for the generation of the next frame. Note that since the ratio of the noise (95%) in the noisy previous frame is much larger than the ratio of the previous frame (5%), the error or complete loss of the previous frame (earlier frames can be used) has almost no impact on the generation of subsequent frames.
Real-time video generation. The frame rate of video playback at the receiver depends on Promptus’s image generation speed, making real-time generation crucial. Promptus’s image generation consists of five parts: prompt dequantization, prompt composition, prompt interpolation, adding noise to previous frame and Stable Diffusion image generation. The first four parts only involve simple linear calculations, so their time consumption can be ignored. The speed of Stable Diffusion becomes the bottleneck. To address this, we adopt the StreamDiffusion [28], which accelerates Stable Diffusion image generation to 100 FPS through batch processing. By this, Promptus achieves real-time video generation on the receiver.
5 Evaluation
Our four main evaluation results are as follows: First, Promptus achieves scalable bitrate, and its quality advantage over the baselines becomes more significant as the target bitrate decreases. Second, Promptus is general to different video domains, and the more complex the video content, the greater the quality advantage of Promptus compared to the baselines. Third, under real-world network traces, Promptus enhances the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Fourth, Promptus can generate videos from prompts in real-time at a speed exceeding 150 FPS.
5.1 Experiment Setup
Test videos: To validate the generalizability of Promptus across different video domains, we selected 7 videos from 4 datasets with vastly different content, as summarized in Table 1. Specifically, the domains of these videos span natural landscapes and human activities, outdoor long-range scenes and indoor close-up scenes, real-world scenes and CG-synthesized scenes, 3D gaming scenes and 2D animations. All videos are cropped and resized to a resolution of 512*512, with a frame rate of 30 FPS. Since Promptus uses a pre-trained and frozen Stable Diffusion, it does not involve model training and does not require training videos. However, for fairness, we still checked that the above test videos were not used during the pre-training of Stable Diffusion.
Table 1: Test videos summary
Dataset #Videos #frames Description
QST [53] 2 300 Natural landscapes, outdoor distant view
UVG [32] 2 300 Human activity, face, hand, indoor close-up
GTA-IM [17] 2 300 3D Game recording, CG-synthesized scenes
Animerun [43] 1 60 2D animation, cartoon
Total 7 960
Baselines: We compare Promptus with two baselines: H.265 [5] and VAE [27, 38]. H.265 is an advanced traditional codec that achieves compression by utilizing hand-designed intra-frame prediction, motion compensation, and transform coding techniques. Moreover, H.265 can control the bitrate of the encoded video by adjusting the quantization parameters, balancing quality and bitrate. Since it does not involve learnable parameters, it has good generality. VAE is a deep learning-based neural codec. Both its encoder and decoder are trainable neural networks. The encoder maps the input image to a low-dimensional latent variable, while the decoder reconstructs the image from the latent variable. The dimensionality of the latent variable is usually much smaller than that of the original image, thus achieving compression. For inter-frame compression of VAE, we first quantize the latent variables of each frame and then encode them into a video using H.265. The quantization process causes almost no quality degradation, and the quality degradation is mainly due to video encoding. Therefore, we adjust the target bitrate of video encoding to balance the quality and bitrate of VAE. For fairness, the training set of VAE is the same as that of Stable Diffusion, both consisting of 5.85 billion images [6].
Metric: To evaluate video quality, we adopt the LPIPS [55] (Learned Perceptual Image Patch Similarity) instead of the traditional SSIM and PSNR. This is because LPIPS has a significantly higher correlation with human subjective ratings compared to SSIM and PSNR [51, 55], which means that LPIPS better reflects human subjective perception of video quality. LPIPS uses a pre-trained neural network to extract layer-wise features from the original image and the distorted image, and then calculates the normalized L2 distance between the features as the LPIPS value. Therefore, a smaller LPIPS value indicates a higher quality of the distorted image. Specifically, we use VGG [41] pre-trained on ImageNet [18] as the feature extractor for LPIPS.
Network trace: We collected 6 network traces from real-world scenarios such as subways, driving, and walking, under 2G, 3G, and 4G networks. The traces consisted of 1 from a 4G network, 2 from 3G, and 3 from 2G. Each trace lasted between 5 and 30 seconds, totaling over 100 seconds. Our traces aimed to test various weak network conditions such as poor signal coverage, network overload, high-speed movement, and frequent switching of mobile communication networks. The average bandwidth per second ranged between 50 kbps and 4000 kbps, while the one-way network delay was approximately 30 ms to 100 ms.
Testbed: To test the streaming performance, we use WebRTC [1] to deploy Promptus and the baselines. WebRTC is a real-time video communication framework that dynamically probes the currently available bandwidth to avoid congestion. To simulate real network conditions, we use Mahimahi [34] to replay the above network traces. The queue length of Mahimahi is set to 60, and the drop-tail strategy is adopted.
5.2 Trade off between bitrate and quality
This section demonstrates how different parameter configurations affect the quality-bitrate tradeoff of Promptus.
Figure 7: Visualization of the fitting results at different ranks. It can be seen that as the rank increases, the prompt can fit more details. When the rank is 4, the earrings are lost. When the rank increases to 16, the earrings are successfully fitted.
Prompt rank. We illustrate the variation in visual quality of Promptus under different ranks, as depicted in Figure 8. It indicates that the higher the prompt rank, the higher the quality of Promptus. For example, when the prompt rank increases from 4 to 16, the LPIPS decreases from 0.265 to 0.221 (on the line with a keyframe interval of 1). This is because the larger the rank, the stronger the representational capability of the prompt, allowing it to fit the target image more accurately, as described in §3.2. We present a visualization of the fitting results at different ranks, as shown in Figure 7. It can be seen that as the rank increases, the prompt can fit more details. For instance, when the rank is 4, the earrings are lost in the fitting result, while when the rank increases to 16, the earrings are successfully fitted.
Keyframe interval. Figure 8 also shows the impact of different keyframe intervals on quality. The smaller the keyframe interval, the higher the quality of Promptus. For example, when the prompt rank is 16, reducing the keyframe interval from 10 to 1 decreases the LPIPS from 0.274 to 0.221. This is because, when the interval is smaller, the distance between keyframes in the prompt space is smaller. Since we constrain the prompt space to be temporally smooth, the linear interpolation of keyframe prompts can more accurately approximate the prompts of intermediate frames, as described in §3.3.
Figure 8: The impact of prompt rank and keyframe interval on visual quality. It indicates that visual quality improves with increasing prompt rank and decreasing keyframe interval.
Quality-bitrate tradeoff. Increasing the prompt rank and reducing the keyframe interval lead to a rapid rise in bitrate while improving quality. Therefore, we present the tradeoff between bitrate and quality for Promptus, as shown in Figure 9. First, it illustrates that overall, the quality monotonically increases with the increase in bitrate. Thus, to optimize quality, Promptus sends prompts at a bitrate closest to the available bandwidth. Second, Promptus can achieve scalable bitrates. For example, by adjusting the rank in the range of 4 to 32 and the keyframe interval in the range of 2 to 8, Promptus’s bitrate spans from 113 kbps to 4284 kbps. Third, at the same bitrate, the quality varies for different configurations. For instance, when the bitrate is 550 kbps, the configuration with a rank of 8 and a keyframe interval of 4 has an LPIPS of 0.255, while the configuration with a rank of 16 and a keyframe interval of 8 has an LPIPS of 0.264, which is lower in quality than the former. So at the same bitrate, Promptus tends to choose configurations with smaller keyframe intervals for streaming.
Figure 9: The tradeoff between bitrate and quality for Promptus. The quality monotonically increases with the increase in bitrate. Besides, Promptus can achieve scalable bitrates.
5.3 Compression efficiency
This section demonstrates the compression efficiency of Promptus. Figure 10 shows the CDF of the frame quality for Promptus and baselines under 4 bitrate levels. Since a lower LPIPS represents better visual quality, a leftward shift of the curve in the figure represents more high-quality frames, and thus higher compression efficiency. We also calculate the average LPIPS for each method, represented by the vertical lines in the figure.
(a) 140 kbps
(b) 280 kbps
(c) 360 kbps
(d) 540 kbps
Figure 10: Frame quality CDF and mean of Promptus and baselines at four target bitrates. It indicates Promptus achieves better compression efficiency across all bitrate levels. Besides, the lower the bitrate, the greater the advantage of Promptus.
First, Promptus achieves better compression efficiency across all bitrate levels. For example, in Figure 10, the curves of Promptus are all to the left of the baseline curves. Second, the lower the bitrate, the greater the advantage of Promptus. At a high level, as the bitrate decreases from 540 kbps to 140 kbps, the distance between Promptus’s curves and the baselines’ curves gradually widens. At a low level, when the bitrate is 540 kbps, the average LPIPS of Promptus is 0.018 and 0.085 lower than VAE and H.265, respectively. When the bitrate is 140 kbps, this difference further increases to 0.139 and 0.118. This is because when the bitrate is reduced, H.265 uses coarser quantization and loses many high-frequency details, resulting in blurriness and block artifacts in the video, which significantly impairs the perceptual quality. VAE mitigates this phenomenon by mapping images to a low-dimensional latent space, reserving more bitrate for video coding. Therefore, at most bitrates, the quality of VAE is superior to H.265. However, at extremely low bitrates (such as 140 kbps), the latent space inevitably introduces distortions caused by video coding. At this point, the VAE Decoder introduces a large number of errors when reconstructing the images, causing a significant decrease in VAE’s quality. On the other hand, when the bitrate is reduced, Promptus reduces the representational capability of the prompt rather than degrading the video quality. This prevents Promptus from accurately describing the video content, resulting in slight misalignments in the generated frames. However, thanks to the inherent image generation capability of Stable Diffusion, Promptus’s frames still have good sharpness and details, and thus the perceptual quality is better than VAE and H.265.
5.4 Generality
This section demonstrates the generality of Promptus across different domains. Figure 12 shows the mean of the frame quality for Promptus and two baselines on four different datasets (as described in Table 1).
Figure 11: Visualization of the compression results on different datasets. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Figure 12: Mean frame quality of Promptus and baselines on 4 different datasets. It demonstrates the generality of Promptus across different domains. Besides, the more high-frequency details a video has, the greater the advantage of Promptus.
First, Promptus achieves better compression efficiency on each dataset. This is because, as shown in Figure 12, Promptus achieves lower average LPIPS compared to the baselines on each dataset. To intuitively demonstrate this improvement, we also visualize the compression results of the three methods on the four datasets at lower bitrates (such as 225 kbps), as shown in Figure 11. It can be observed that, compared to the baselines which exhibit blurriness and blocking artifacts caused by compression, Promptus preserves more high-frequency details, resulting in higher perceptual quality.
Second, the more high-frequency details a video has, the greater the advantage of Promptus. For example, for the Animerun dataset with fewer details, the LPIPS of Promptus is 0.015 lower than H.265, which is not a significant advantage. However, for the detail-rich UVG, this difference further expands to 0.121. This is because for 2D animations with large areas of solid colors and simple details, H.265’s intra-frame prediction, block partitioning, and motion compensation techniques can handle them well, so Promptus’s performance gain is small. For detail-rich real-world videos, H.265 discards more high-frequency information during compression, thus damaging the perceptual quality (especially at lower bitrates). Although Promptus also loses high-frequency information from the original image, Stable Diffusion completes the lost high-frequency information based on prior knowledge during generation, resulting in a smaller perceptual quality loss.
5.5 Performance on real-world traces
This section demonstrates the performance of Promptus under real network traces. We use the traces and testbed described in §5.1 to run Promptus and two baselines. Figure 13 shows the CDF and mean of the frame quality. Since the total length of the traces is greater than the total length of the test videos, we loop the test videos to run through the entire traces.
Figure 13: The performance of Promptus under real network traces. On one hand, Promptus’s quality is overall higher than the baselines. On the other hand, Promptus can significantly reduce the ratio of severely distorted frames.
First, Promptus’s quality is overall higher than the baselines. For example, the mean LPIPS of Promptus is 0.111 and 0.092 lower than VAE and H.265, respectively. Second, Promptus can significantly reduce the ratio of severely distorted frames. For instance, only 5.2% of frames in Promptus have LPIPS higher than 0.32, while VAE and H.265 have 94.5% and 96.9%, respectively. These improvements are partly due to Promptus’s excellent compression efficiency, enabling it to provide higher perceptual quality at the same bitrate. On the other hand, since Promptus sends raw prompts without encoding (such as entropy coding), it can precisely control the target, thus making full use of bandwidth.
5.6 Overhead
In this section, we analyze the overhead of Promptus’s generation. We conduct tests on an Nvidia 4090D GPU, using CUDA to accelerate Promptus. The resolution of the generated video is 512*512, and the rank of the prompt is 8.
Table 2: Overhead of generating a frame
Steps Time (ms) Memory (MB)
Prompt dequantization 0.016 -
Prompt composition 0.025 -
Prompt interpolation 0.013 -
Noised previous frame 0.012 -
Stable Diffusion generation 6.160 -
Total 6.226 8952
Table 2 shows the fine-grained overhead of each step in Promptus’s image generation. Specifically, from receiving a prompt to generating a frame, Promptus includes the following steps: prompt dequantization, prompt composition, prompt interpolation, adding noise to the previous frame, and Stable Diffusion image generation. Among them, most steps only involve simple linear computations, so the time overhead is almost negligible. In contrast, Stable Diffusion image generation accounts for the vast majority of the time overhead. Therefore, we adopt StreamDiffusion [28], which accelerates Stable Diffusion to real-time through batch processing techniques and implementation optimizations. The difference is that compared to the original StreamDiffusion, Promptus removes the CLIP module that converts text to embeddings, thus further accelerating image generation to 6.160 ms. In summary, the total time overhead of Promptus’s image generation is 6.226 ms, achieving real-time.
6 Limitation
The time overhead of prompt fitting. Although Promptus can achieve real-time video generation at the receiver, prompt fitting cannot be performed in real-time at the sender. This is because prompt fitting requires iterative gradient descent. Although a single iteration is fast (on the same order of magnitude as the time overhead of real-time generation), the total time overhead is high due to the large number of iterations required for convergence (such as 500 iterations). To address this, using more efficient gradient descent algorithms to reduce the number of iterations required for convergence is expected to accelerate prompt fitting to real-time.
The latency of prompt interpolation. At the receiver side, Promptus obtains the prompts of intermediate frames through prompt interpolation of keyframes. This means that if an intermediate frame needs to be generated and played, it is necessary to wait until the subsequent keyframe is received, which introduces additional latency. To address this issue, designing keyframe extrapolation algorithms to replace interpolation is a future research direction.
Non-uniform keyframes. According to §4.1, Promptus sends keyframes uniformly based on the keyframe interval. However, different segments of a video often have different rates of change. For rapidly changing segments, the distance between keyframes in the prompt space is large, resulting in a poor approximation of intermediate frames using linear interpolation. A solution is to reduce the keyframe interval and send keyframes more densely. However, there also exist smoothly changing segments in the video, where densely sending keyframes brings little improvement to quality, thus wasting bandwidth. In summary, sending keyframes uniformly in this paper is inefficient. In the future, adaptive keyframe needs to be designed to transmit densely in rapidly changing segments and sparsely in smoothly changing segments.
7 Conclusion
In this paper, we propose Promptus, a novel system that replaces video streaming with prompt streaming by inverting video frames into prompts for Stable Diffusion. To ensure pixel alignment, a gradient descent-based prompt fitting framework is proposed. To achieve adaptive bitrate for prompts, a low-rank decomposition-based bitrate control algorithm is introduced. For inter-frame compression of prompts, a temporal smoothing-based prompt interpolation algorithm is proposed. Evaluations across various video domains and real network traces demonstrate Promptus can enhance the perceptual quality by 0.111 and 0.092 (in LPIPS) compared to VAE and H.265, respectively, and decreases the ratio of severely distorted frames by 89.3% and 91.7%. Moreover, Promptus achieves real-time video generation from prompts at over 150 FPS. By pioneering the replacement of video codecs with prompt inversion and introducing prompt streaming, Promptus revolutionizes efficient video communication, surpassing the limitations of traditional approaches and paving the way for a new era of video streaming.